Meg is 20 years younger than Phillip. In three years Phillip will be twice as old as Meg will be. How old is each now?
Answer by Edwin McCravy(20056)


Meg is 20 years younger than Phillip. In three years Phillip will be twice as old as Meg will be. How old is each now?
`
>>...Meg is 20 years younger than Phillip...<<
`
So M = P - 20
`
>>...In three years Phillip will be...<<<
`
In 3 years Philip will be P + 3
`
>>...twice as old as Meg will be...<<
`
In 3 years Meg will be M + 3
`
So
`
P + 3 = 2(M + 3)
`
So we have the system of two equations and 2 unknowns:
`
M = P - 20
P + 3 = 2(M + 3)
`
Use substitution to solve this system.
`
Answer: M = 17, P = 37.
`
Checking:
`
Meg is 17, which is 20 years younger than Phillip, who is 37. In three years Phillip will be 40, and then he will be twice as old as Meg will be, for she will then be be 20.
`
Edwin